OpenStudy (anonymous):

Suppose G(x) =2x^2+x+4. Find a number b such that G'(b)=7.

OpenStudy (xishem):

First, you need to differentiate G(x) with respect to x: G'(x) = 4x+1 Then, you need to solve it for some input b which outputs 7: 7 = 4x+1 6 = 4x x = 6/4 = 3/2 = 1.5 b = 1.5